OpenStudy (anonymous):

8. (6pts).You have an urn that contains 12 balls, 5 yellow, 4 red, and 3 green. You randomly select two balls. a) What is the probability that you select two balls of the same color?

OpenStudy (anonymous):

compute for each, then add

OpenStudy (anonymous):

for example two yellow is \[\frac{5\times 4}{12\times 11}\]

OpenStudy (anonymous):

you go the rest?

OpenStudy (anonymous):

Are you sure you can add the probabilities of dependent events?

OpenStudy (anonymous):

i am not sure what you mean by "dependent" you can add them because they are "disjoint" two yellow or two red or two green

OpenStudy (anonymous):

oh I understand. Thank you